OpenStudy (anonymous):

15fc f-c ----- - ---- subtract and simplify by f^2-c^2 f+c removing a factor of 1 if possible

OpenStudy (anonymous):

i think the simplified answer for this is \[-( f ^{2}-17fc+c ^{2} )/ (f+c)(f-c)\]

OpenStudy (anonymous):

\[(15fc-(f-c)^2)/(f^2-c^2)\] \[=(15fc-(f^2-2fc+c^2))/(f^2-c^2)\] \[=(15fc-f^2+2fc-c^2)/(f^2-c^2)\] \[=(13fc-f^2-c^2)/(f^2-c^2)\]